Scott wants to swim across a river that is 400 meters wide. He begins swimming perpendicular to the shore he started from but ends up 100 down river from where he started because of the current. How far did he actually swim from his starting point?
Answer by fractalier(6550)

Assuming that his travel was in a straight line, which it probably wasn't, and assuming that the current was constant throughout his swim, which it wasn't, the problem becomes a simple Pythagorean theorem problem.
You need to find the hypotenuse of a right triangle 400 m on one leg and 100 m on the other.
So c^2 = a^2 + b^2 = 400^2 + 100^2 = 170000
And c = square root (170000) or about 412.3 m